If fixed costs are $561,000. and the unit contribution margin is $8.00, what is the break-even point in units if the variable costs are increased by $.50 per unit?

can someone please add the calculation or formula used.
Answer #1

If variable costs increase by 0.50 per unit then revised contribution margin = 8 - 0.5 = $7.5 per unit

Break even point (units) = Fixed costs / contribution margin = 561000/7.5 = 74800 units